Past history shows that 45% of university students prefer Brand C cola. A random sample of 10 students is selected.   

What is the probability that exactly 3 of the students prefer Brand C cola?
Answer #1

Solution :

Given that ,

p = 0.45

1 - p = 1 -0.45 = 0.55  

n = 10

Using binomial probability formula ,

P(X = x) = ((n! / (n - x)!) * px * (1 - p)n - x

P(X = 3) = ((10! / (10- 3)!) * 0.453 * 0.5510 - 3

= 0.1665

Probability = 0.1665